# Implicit Neural Representations for Spatial Transcriptomics

### v1

@@ 6 Background
Spatial Transcriptomics (ST) is a technique that representing the gene expression at individual spots within a slice of sample, providing a detailed picture of cellular activity and composition in the micro-environment. ST sequencing is typically performed on fresh frozen tissue samples and is broadly categorized into two main types: sequencing-based ST and imaging-based ST. Despite differences in methodology, both technologies deliver complementary measurements of in situ gene expression. The sequencing-based ST technologies, including Stereo-seq, 10x Visium, Slide-seqV2, usually rely on spatially indexed surfaces to encode spatial information (Supplementary Fig. 7). These methods employ different spatial indexing strategies: 10x Visium uses a microarrayer spotting robot, these spots are 50–200 $\mu m$ in size, to deliver a unique barcode to a fixed, known location on the surface of a slide [51]; Slide-seq applies 10 $\mu m$ diameter nanobeads for spatial barcoding [51]; Stereo-seq uses DNA nanoball-patterned arrays for in situ RNA capture to create high-resolution spatial transcriptomics sequencing [2]. Following barcoding, mRNA diffuses to spatially indexed primers, undergoes reverse transcription, cDNA amplification, and short-read sequencing. This process generates transcript reads paired with barcode sequences, localizing each detected transcript to specific pixels.
@@
7 Model Details & Insights
@@
As mentioned in the main body, it is observed that when applied to ST with low spatial density, SIREN [43] outperforms FFN [48] as the backbone of SUICA, while for high-resolution ST slices, FFN prevails. To better showcase such phenomenon, we provide two representative cases of the mouse embryos E9.5 and E16.5 in Supplementary Tab. 5. We globally set the angular velocity $\omega$ of the periodical activation function in SIREN [43] as 30 which is adopted by default and the mapping size of the Random Fourier Features as 256. For more information for these two magic numbers, please refer to the respective papers. Note that we do not intend to showcase any superiority among INR variants. Since we adopt a modular design that is agnostic to specific INR implementations, we are able to take advantage of different implementations to meet varying requirements.
@@
FFN [48] 0.55 0.752 0.799 1.47 0.807 0.761
SIREN [43] 0.52 0.766 0.814 1.62 0.786 0.745
Table 5: Quantitative results of SUICA empowered by different INR backbones.
@@
Recently, researchers are employing explicit or hybrid alternatives [32, 3] to replace the fully implicit representations [48, 43, 23] for a faster inference speed. For a more comprehensive benchmarking, we here compare SUICA with the trainable hash-based encoding [32] (denoted as NGP) on Stereo-seq MOSTA dataset [2]. The quantitative results are shown in Supplementary Tab. 6. For the implementation, we re-compile the library of tiny-cuda-nn to support float32, to align with other methods.
@@
Table 6: Quantitative comparison on Stereo-seq MOSTA dataset [2] between SUICA and NGP [32].
@@
It is easy to notice that the network architecture in SUICA is almost fully MLP-based except for the GCN. Such architecture does not explicitly involve any local inductive bias that CNNs usually do. The underlying reason for doing so is that along the sequencing depth of ST, it is assumed that the channels (i.e., specific gene segments) are permutation-invariant even though there might be some statistical correlations. In this way, it is not supposed to make use of the context information along the sequencing depth but for the spatial adjacency, we model such context by the GCN.
@@
Note that by aid of GAE, we reduce the dimension of raw gene expressions from thousands to 32, which is recovered later by the decoder. The job of the GAE can be replaced by PCA (Principal Component Analysis), by a pre-fitted linear projection. However, it is obviously challenging for the linear transformation of PCA and inverse PCA to model the complex distribution of ST, leading to a great loss in the reconstructed signal. Empirically, it is observed that when encoding and decoding with PCA and inverse PCA, the IoU between the predicted zero maps and the ground truth decreases to 0, indicating a complete failure of recovering the sparsity, while SUICA maintains an IoU of around 0.9.
@@
Variational Autoencoders (VAEs) are a common alternative to AEs, designed to generate samples from a probabilistic latent space, typically modeled as Gaussian distributions. While VAEs offer generative capabilities through variational inference and learn smoother, more continuous latent spaces compared to AEs, they often produce blurry outputs [60] and struggle to match AEs in reconstruction accuracy [7]. These issues might lead to ambiguity in gene expression that is not wanted. These limitations stem from the inherent trade-off between the quality of latent representation learning and reconstruction fidelity in VAEs. Although variants like beta-VAE [13] allow for adjustable weighting between these objectives, the issue persists. In our setting, where numerical fidelity and reconstruction accuracy take precedence over generative and latent representation quality, we select AE over VAE for SUICA’s design, leveraging its ability to accurately reconstruct high-dimensional raw gene expression data.
@@
8 STAGE
In the main text, we involve STAGE [21] for both quantitative and qualitative comparisons as a baseline for histology-free super-resolution of gene expressions. In addition to solely showing the superiority of SUICA, we would like to shed some light on this matter.
Regarding the benchmarking task of super-resolution, we identify two critical sub-tasks, namely the reconstruction of gene expression and the spatial mapping. STAGE models the super-resolution as a generative task, where an AE is adopted to self-regress the raw representations while enforcing the latent as the corresponding 2D coordinates. In this way, the two sub-tasks are coupled as one. Considering the challenges posed, the closely coupled paradigm may not lead to satisfactory results. On the contrary in SUICA, the two sub-tasks are well decoupled, where the spatial mapping is performed by the INR while the description as well as the reconstruction is taken care of the GAE. The progressive training paradigm also guarantees that each module is doing the assigned job, keeping the coupling at minimum.
@@ 8 STAGE
9 MERFISH
MERFISH is an imaging-based ST technique that enables the highly multiplexed imaging of RNA molecules in cells while maintaining their spatial context. Compared to the Slide-seqV2, 10x Visium, and Stereo-seq technologies used in our manuscript, MERFISH data can quantify significantly fewer genes for individual cells. In Supplementary Tab. 7, we evaluate the performance of SUICA using a human heart MERFISH dataset [10] with 228,635 cells and 238 genes. SUICA achieves a significant lower mean absolute error (MAE) and mean square error (MSE), and the cosine similarity is at least 2.28% higher than other methods, showing its high numerical fidelity. Following the bio-conservation analysis scheme, we annotate the cell types for predicted spots of MERFISH data using dominant cell types in individual clusters. We visualize the spatial bio-conservation by coloring the cell types in Supplementary Fig. 8. The results show that the SUICA predictions mimic the ground-truth evidenced the capability of SUICA to predict unseen spots in the MERFISH datasets.
Figure 8: Spatially visualized predicted spots on MERFISH human heart [10].
@@ 9 MERFISH
Nearest Interp. 6.43 6.53 0.814 0.641 0.476
FFN [48] 6.06 5.86 0.840 0.717 0.554
SIREN [43] 5.54 4.98 0.864 0.759 0.606
STAGE [21] 5.48 5.10 0.870 0.709 0.558
@@
Table 7: Quantitative results on MERFISH human heart [10].
10 Data Pre-processing
@@
Note that the outputs of conventional INR-based tasks are strictly bounded, where the linear layer is usually clamped or compressed by sigmoid. For ST data, we do not have such assumption and only employ ReLU as the final activation. Therefore, as a common practice in ST analysis [28, 17], we remove the genes whose expressions are overly high and normalize each cell by total counts over all genes of the cell so that every cell has the same total count after normalization while keeping the original sparsity.
11 Limitations
One of the significant limitations in SUICA roots in the case-by-case training paradigm of INRs. Similar to NeRF [29], for each case, i.e., each ST slice, we need to training a new model from scratch. Considering the overwhelming heterogeneity of different ST data, we temporarily compromise in this issue but incorporate domain knowledge to make SUICA generalizable is an interesting future work.
Another potential limitation is that SUICA prefers high-quality ST data, in terms of both spatial density and sequencing depth, as can be indicated from the experimental results we provide. SUICA exhibits the most performance gain in Stereo-seq compared to other platforms and when the gene expressions no longer meet the super-high dimensional sparse assumption, SUICA will degenerate to a vanilla INR, as is consistent with the results of MERFISH in Supplementary Tab. 7.

### v2

@@ Appendix A Background
Spatial Transcriptomics (ST) is a technique that representing the gene expression at individual spots within a slice of sample, providing a detailed picture of cellular activity and composition in the micro-environment. ST sequencing is typically performed on fresh frozen tissue samples and is broadly categorized into two main types: sequencing-based ST and imaging-based ST. Despite differences in methodology, both technologies deliver complementary measurements of in situ gene expression. The sequencing-based ST technologies, including Stereo-seq, 10x Visium, Slide-seqV2, usually rely on spatially indexed surfaces to encode spatial information (Figure 6). These methods employ different spatial indexing strategies: 10x Visium uses a microarrayer spotting robot, these spots are 50–200 $\mu m$ in size, to deliver a unique barcode to a fixed, known location on the surface of a slide (Tian et al., 2023); Slide-seq applies 10 $\mu m$ diameter nanobeads for spatial barcoding (Tian et al., 2023); Stereo-seq uses DNA nanoball-patterned arrays for in situ RNA capture to create high-resolution spatial transcriptomics sequencing (Chen et al., 2022a). Following barcoding, mRNA diffuses to spatially indexed primers, undergoes reverse transcription, cDNA amplification, and short-read sequencing. This process generates transcript reads paired with barcode sequences, localizing each detected transcript to specific pixels.
@@
We also want to mention the fact that ST data are not always delivered jointly with histology images as a reference. For example, Slide-seq, DBiT-seq, MERFISH offer high-resolution spatial gene expression data but lack corresponding histological images. Even in cases where histology is available, challenges in precise alignment between histological features and ST data can arise due to tissue deformation, sectioning artifacts, or registration inaccuracies. These limitations highlight the importance of designing reference-free models that do not rely on histological guidance, making them more broadly applicable and robust across different ST platforms and experimental conditions.
Appendix B Model Details & Insights
@@
As mentioned in the main body, it is observed that when applied to ST with low spatial density, SIREN (Sitzmann et al., 2020) outperforms FFN (Tancik et al., 2020) as the backbone of SUICA, while for high-resolution ST slices, FFN prevails. To better showcase such phenomenon, we provide two representative cases of the mouse embryos E9.5 and E16.5 in Table 4. We globally set the angular velocity $\omega$ of the periodical activation function in SIREN (Sitzmann et al., 2020) as 30 which is adopted by default and the mapping size of the Random Fourier Features as 256. For more information for these two magic numbers, please refer to the respective papers. Note that we do not intend to showcase any superiority among INR variants. Since we adopt a modular design that is agnostic to specific INR implementations, we are able to take advantage of different implementations to meet varying requirements.
Table 4: Quantitative results of SUICA empowered by different INR backbones. MSE: ×$\times 10^{-2}$.
@@
FFN (Tancik et al., 2020) 0.55 0.752 0.799 1.47 0.807 0.761
SIREN (Sitzmann et al., 2020) 0.52 0.766 0.814 1.62 0.786 0.745
@@
Recently, researchers are employing explicit or hybrid alternatives (Müller et al., 2022; Chen et al., 2022b) to replace the fully implicit representations (Tancik et al., 2020; Sitzmann et al., 2020; Lindell et al., 2022) for a faster inference speed. For a more comprehensive benchmarking, we here compare SUICA with the trainable hash-based encoding (Müller et al., 2022) (denoted as NGP) on Stereo-seq MOSTA dataset (Chen et al., 2022a). The quantitative results are shown in Table 5. For the implementation, we re-compile the library of tiny-cuda-nn to support float32, to align with other methods.
Table 5: Quantitative comparison on Stereo-seq MOSTA dataset (Chen et al., 2022a) between SUICA and NGP (Müller et al., 2022). MAE/MSE: ×$\times 10^{-2}$.
@@
It is easy to notice that the network architecture in SUICA is almost fully MLP-based except for the GCN. Such architecture does not explicitly involve any local inductive bias that CNNs usually do. The underlying reason for doing so is that along the sequencing depth of ST, it is assumed that the channels (i.e., specific gene segments) are permutation-invariant even though there might be some statistical correlations. In this way, it is not supposed to rely on the context information along the sequencing depth, but for the correlation among cells, we model such context by the GCN.
@@
In SUICA, we reduce the dimension of raw gene expressions from thousands to 32 by the encoder of GAE, which is recovered later by the decoder. The function of the GAE could be replaced by PCA (Principal Component Analysis), by a pre-fitted linear projection. However, it is obviously challenging for the linear transformation of PCA and inverse PCA to model the complex distribution of ST, leading to a great loss in the reconstructed signal. Empirically, it is observed that when encoding and decoding with PCA and inverse PCA, the IoU between the predicted zero maps and the ground truth decreases to 0, indicating a complete failure of recovering the sparsity, while SUICA maintains an IoU of around 0.9.
@@
Variational Autoencoders (VAEs) are a common alternative to AEs, designed to generate samples from a probabilistic latent space, typically modeled as Gaussian distributions. While VAEs offer generative capabilities through variational inference and learn smoother, more continuous latent spaces compared to AEs, they often produce blurry outputs (Zhao et al., 2017) and struggle to match AEs in reconstruction accuracy (Dai et al., 2020). These issues might lead to ambiguity in gene expression that is not wanted. These limitations stem from the inherent trade-off between the quality of latent representation learning and reconstruction fidelity in VAEs. Although variants like beta-VAE (Higgins et al., 2017) allow for adjustable weighting between these objectives, the issue persists. In our setting, where numerical fidelity and reconstruction accuracy take precedence over generative and latent representation quality, we select AE over VAE for SUICA’s design, leveraging its ability to accurately reconstruct high-dimensional raw gene expression data.
@@
Table 6: Quantitative benchmarking results of spatial imputation on two Visium ST cases (Palla et al., 2022; Wei et al., 2022). Note that for Visium-Human Brain, there is no annotation of cell type for the evaluation of ARI, while reference ARI is 0.428 for Mouse Brain. Bold figures are best scores and underlined figures are second-best. MAE/MSE: ×$\times 10^{-2}$ for Human Brain while ×$\times 10^{-1}$ for Mouse Brain.
Methods Visium-Human Brain Visium-Mouse Brain
[…]
FFN (Tancik et al., 2020) 5.76 0.881 0.772 0.786 0.402 5.95 5.85 0.832 0.741 0.581 0.000587
SIREN (Sitzmann et al., 2020) 6.58 0.933 0.756 0.747 0.196 5.35 4.29 0.878 0.804 0.647 0.359
STAGE (Li et al., 2024) 6.19 0.805 0.795 0.772 0.223 4.55 3.20 0.918 0.825 0.666 0.140
TRIPLEX (Chung et al., 2024) 4.75 0.560 0.881 0.850 0.319 9.35 14.0 0.00 -0.00682 -0.00715 0.358
UNIv2 (Chen et al., 2024) 7.30 1.41 0.723 0.633 0.129 6.94 7.88 0.790 0.631 0.425 0.228
SUICA (Ours) 4.99 0.567 0.860 0.846 0.445 3.68 2.45 0.932 0.800 0.660 0.393
Appendix C STAGE
In the main text, we involve STAGE (Li et al., 2024) for both quantitative and qualitative comparisons as a baseline for histology-free spatial imputation (densification) of gene expressions. In addition to solely showing the superiority of SUICA, we would like to shed some light on this matter.
Regarding the benchmarking task of imputation, we identify two critical sub-tasks, namely the reconstruction of gene expression and the spatial mapping. STAGE models imputation as a generative task, where an AE is adopted to self-regress the raw representations while enforcing the latent as the corresponding 2D coordinates. In this way, the two sub-tasks are coupled as one. Considering the challenges posed, the closely coupled paradigm may not lead to satisfactory results. On the contrary in SUICA, the two sub-tasks are well decoupled, where the spatial mapping is performed by the INR while the description as well as the reconstruction is taken care of the GAE. The progressive training paradigm also guarantees that each module is doing the assigned job, keeping the coupling at minimum.
@@ Appendix C STAGE
Appendix D Human and Mouse Brains Visium
Different from Stereo-seq datasets, Visium is a lower-resolution but more affordable ST technology, on which we additionally compare the model performance. The experiments are conducted under the task of spatial imputation. Due to the lack of independent cell-type annotations, ARI is not applicable in Visium-Human Brain dataset (Wei et al., 2022). On the Visium-Mouse Brain dataset (Palla et al., 2022), SUICA demonstrates improved ARI, which indicates that the predictions can more accurately describe the underlying cellular heterogeneity. Note that for readers’ information, we also involve SOTA histology-aided imputation methods TRIPLEX (Chung et al., 2024) and UNI (Chen et al., 2024) for comparison, which additionally has access to extra reference. We have also found that TRIPLEX is rather sensitive to dataset-specific characteristics.
Appendix E MERFISH
MERFISH is an imaging-based ST technique that enables the highly multiplexed imaging of RNA molecules in cells while maintaining their spatial context. Compared to the Slide-seqV2, 10x Visium, and Stereo-seq technologies used in our manuscript, MERFISH data can quantify significantly fewer genes for individual cells. In Table 7, we evaluate the performance of SUICA using a human heart MERFISH dataset (Farah et al., 2024) with 228,635 cells and 238 genes, with the setting of spatial imputation. SUICA achieves a significant lower mean absolute error (MAE) and mean square error (MSE), and the cosine similarity is at least 2.28% higher than other methods, showing its high numerical fidelity. Following the bio-conservation analysis scheme, we annotate the cell types for predicted spots of MERFISH data using dominant cell types in individual clusters. We visualize the spatial bio-conservation by coloring the cell types in Figure 7. The results show that the SUICA predictions mimic the ground-truth evidenced the capability of SUICA to predict unseen spots in the MERFISH datasets.
Figure 7: Spatially visualized predicted spots on MERFISH human heart (Farah et al., 2024).
Table 7: Quantitative results of spatial imputation on MERFISH human heart (Farah et al., 2024). MAE/MSE: ×$\times 10^{-1}$.
@@ Appendix E MERFISH
FFN (Tancik et al., 2020) 6.06 5.86 0.840 0.717 0.554
SIREN (Sitzmann et al., 2020) 5.54 4.98 0.864 0.759 0.606
STAGE (Li et al., 2024) 5.48 5.10 0.870 0.709 0.558
@@
Appendix F Ablation on Resolution
To help with the data-efficiency analysis of SUICA, we perform experiments under the task of spatial imputation with E16.5 embryo of Stereo-seq MOSTA (Chen et al., 2022a), the results of which are shown in Table 8. The percentage refers to the the proportion of training samples with regard to the whole dataset, while the test set remains the same (20% of all spots).
Table 8: Ablation study on the data-efficiency of SUICA. % represents the proportion of the spots used for training while test set remains 20%. MAE/MSE: ×$\times 10^{-2}$.
[…]
80% 8.01 1.47 0.807 0.761
60% 7.96 1.52 0.801 0.752
40% 8.00 1.59 0.790 0.739
20% 8.14 1.62 0.786 0.738
Appendix G Data Pre-processing
@@
Note that the outputs of conventional INR-based tasks are strictly bounded, where the linear layer is usually clamped or compressed by sigmoid. For ST data, we do not have such assumption and only employ ReLU as the final activation. Therefore, as a common practice in ST analysis (Melsted et al., 2021; Hwang et al., 2018), we remove the genes whose expressions are overly high and normalize each cell by total counts over all genes of the cell so that every cell has the same total count after normalization while keeping the original sparsity.
Appendix H Limitations
One of the significant limitations in SUICA roots in the case-by-case training paradigm of INRs. Similar to NeRF (Mildenhall et al., 2020), for each case, i.e., each ST slice, we need to training a new model from scratch. Considering the overwhelming heterogeneity of different ST data, we temporarily compromise in this issue but acknowledge that incorporating domain knowledge to make SUICA generalizable is an interesting future work.
Another potential limitation is that SUICA prefers high-quality ST data, in terms of both spatial density and sequencing depth, as can be indicated from the experimental results we provide. SUICA exhibits the most performance gain in Stereo-seq compared to other platforms and when the gene expressions no longer meet the super-high dimensional sparse assumption, SUICA will degenerate to a vanilla INR, as is consistent with the results of MERFISH in Table 7.